# Efficient Mixture-of-Expert for Video-based Driver State and Physiological Multi-task Estimation in Conditional Autonomous Driving

### v1

@@ Abstract
Road safety remains a critical challenge worldwide, with approximately 1.35 million fatalities annually attributed to traffic accidents, often due to human errors. As we advance towards higher levels of vehicle automation, challenges still exist, as driving with automation can cognitively over-demand drivers if they engage in non-driving-related tasks (NDRTs), or lead to drowsiness if driving was the sole task. This calls for the urgent need for an effective Driver Monitoring System (DMS) that can evaluate cognitive load and drowsiness in SAE Level-2/3 autonomous driving contexts. In this study, we propose a novel multi-task DMS, termed VDMoE, which leverages RGB video input to monitor driver states non-invasively. By utilizing key facial features to minimize computational load and integrating remote Photoplethysmography (rPPG) for physiological insights, our approach enhances detection accuracy while maintaining efficiency. Additionally, we optimize the Mixture-of-Experts (MoE) framework to accommodate multi-modal inputs and improve performance across different tasks. A novel prior-inclusive regularization method is introduced to align model outputs with statistical priors, thus accelerating convergence and mitigating overfitting risks. We validate our method with the creation of a new dataset (MCDD), which comprises RGB video and physiological indicators from 42 participants, and two public datasets. Our findings demonstrate the effectiveness of VDMoE in monitoring driver states, contributing to safer autonomous driving systems. The code and data will be released.
@@ I Introduction
Road safety is still an extreme challenge for societies [1, 2]. The World Health Organization (WHO) reported that approximately 1.35 million people die each year as a result of road traffic accidents, with human error being a significant contributing factor. While autonomous technologies promise to reduce human error—the leading cause of traffic accidents—there remains a critical period where human drivers must share control with automated systems. Particularly, at the Society of Automotive Engineers (SAE) Level-2 (L2) automation [3], the vehicle can control both steering and acceleration/deceleration but still requires the driver to remain actively engaged and ready to take over at any moment; while with SAE Level-3 (L3) automation, the vehicle can handle most driving tasks but still requires human drivers to step in when prompted by the take-over requests (TORs). Both have shifted the role of the driver from an operator to a supervisor. As a result, drivers are more likely to experience low cognitive load if driving is the sole task, or more likely to be cognitively over-demanding, if drivers are engaged in non-driving-related tasks (NDRTs), both can be detrimental to driving performance [4]. For example, the cognitive overload induced by multi-dimensional cognitive resource consumption can lead to a limited field of vision [5], and a reduced ability to foresee potential hazards [6]. Futher, the effect of long-time driving fatigue and the high cognitive load as a result of NDRTs may co-exist, leading to impaired driving performance [7, 8]. Thus, the Driver Monitoring System (DMS) that is capable of estimating drivers’ cognitive load and drowsiness is still critical to enhancing the safety and reliability of SAE L2/L3 vehicles.
@@
However, most of the previous methods were based on single frame detection, which loses temporal variation in both surface facial expression and internal physiological reflection. Compared to image-based methods, video-based methods [20, 17, 21] can make effective use of continuous temporal features and have the potential to enhance detection accuracy. However, existing methods still have the following shortcomings: (1) some video-based drowsiness detection methods [22] directly utilize continuous frames as the input, leading to significant computational resources and interference from redundant information; (2) some methods [17, 23] tried to replace video by key optimized facial features (e.g., landmarks, eye and mouse areas), while they cannot assess physiological features without supplementing signals from physiological sensors; (3) current video-based DMS focused mostly on single-task fatigue [20, 17, 21] or distraction [24] estimation. However, driver states do not appear independently; instead, the interaction effects among multiple states make driver state estimation complicated. For example, in the context of conditional autonomous driving (i.e., driving with SAE L2/L3 vehicles), the cognitive load can moderate both fatigues [25] and physiological reflection [26] . However, few non-contact solutions were proposed for cognitive load estimation, fatigue detection and physiological reflections. Given that training separate models for each task incurs high deployment costs and reduces iterative efficiency [27], it is urgent to develop a unified model and make full use of the association across drivers’ multiple states.
@@
To address the aforementioned issues, we propose an RGB video-based multi-task DMS (VDMoE). Specifically, based on previous studies [17, 34], we first utilized the key facial features to reduce redundant information and computational costs. At the same time, considering cardiac and respiration activities are crucial predictors of driver states [26, 25], and are also important driver health indicators, we integrated the remote Photoplethysmography (rPPG) technology [35] into our model. Since the facial landmark and local eye and mouth region process contain limited temporal color change of the facial skin pixel points, which can be converted to a change of the blood volume under the skin [27], we first selected key facial regions of interest based on facial landmarks. Then, by transforming the color space (from RGB to YUV) as well as by band-pass denoising, we obtained the alternative multi-modal information input Spatial-temporal Map (STMap). Next, to fully leverage the dependencies between different tasks, the Mixture-of-Experts (MoE) [36] structure is introduced. Being different from the classic MoE structure, the heterogenous gating mechanism and spatio-temporal expert separation are designed for the multi-modal alternative inputs to improve the multi-task performance. Besides, recent video-based DMS used various types of neural networks (e.g., CNN [20], Transformer [17]) that are suitable for processing different types of data or extracting different levels of feature information [37]. However, given the demand for real-time assessment of DMS, networks with excessive parameters and computational complexity will pose a challenge to deployment hardware. Therefore, we instantiated the nonlinear feature learning component of the basic block of VDMoE as a simple two-layer multi-layered perception (MLP) network to replace the previous modules represented by residual convolution [38] or multi-head attention [39]. Lastly, for the optimization goal, to explicitly capture the dependency between different states, we strengthen the learning capacity of VDMoE through one regularization based on prior knowledge from human factors fields [25]. Given there is no proper dataset for multi-task video-based DMS development that considered both the cognitive load and drowsiness in the driving automation context (seeing Table II), a driving simulator experiment with 42 participants was conducted. Extensive experiments based on our dataset illustrate the effectiveness of our method. In all, the contributions of our work are summarized as follows:
@@
  • A multi-task RGB video-based driver state monitoring method (VDMoE) is proposed in this work. This system concentrates on the fatigue and physiological reflection moderated by the multi-dimensional cognitive load in the context of driving automation.
@@
  • To achieve the balance between estimation accuracy and efficiency, we replaced the full video with key facial features (landmarks, eye, and mouse area) as input, and introduced the STMap representing color changes in facial regions to supplement alternative physiological features.
@@
  • Inspired by the superior performance of MoE in multi-task learning, we further optimized the MoE structure to accommodate the multi-modal and spatio-temporal input through heterogeneous gating mechanism and spatio-temporal expert separation. At the same time, we replaced more complex network structures with MLPs for lightweight.
@@
  • A novel prior-inclusive regularization was proposed to enforce the multi-task model’s outputted probability distributions for each state to be aligned with the statistical prior distributions so that we can speed up convergence and reduce the risk of overfitting the model to particular individuals.
@@
  • To validate the effectiveness of our method, the first multi-modal multi-modal cognitive load and drowsiness driving dataset (MCDD) in the L3 context was created. In total, 42 subjects were involved in this dataset, which included about 105,840 second RGB video and cardiac and respiration-related physiological indicators. The cleaned processed dataset will be released.
@@ II-A Physiological Signal-Based Driver State Monitoring
Driver fatigue [40] can negatively affect drivers’ performance when driving with automation, especially in takeover events. Thus, driver state monitoring is necessary to ensure safe driving in the context of driving automation [11]. According to [41, 42], in conditional autonomous vehicles, certain physiological metrics can be strongly correlated with driver state. Therefore, researchers proposed various physiological signal-based driver state monitoring methods. Commonly used psychophysiological measures include electrophysiological signals and human physical movements [43]. For example, Cui et al. [44] designed an interpretable convolutional neural network for drowsiness detection using EEG signals. In addition, due to the difficulty of obtaining EGG signals, methods based on Electrocardiography(ECG) signals [45] and Electrodermal Activity(EDA) signals [46] have also been widely proposed. Yang et al. [34] developed an attention-enabled recognition network with a decision-level fusion architecture to estimate cognitive load based on physiological data. Another study used physiological signals and facial images to detect driver fatigue [20]. However, the above-mentioned studies used physiological data collected through intrusive sensors, i.e., the sensors had to contact with the driver’s body, which makes them infeasible to be used in the actual driving environment.
@@ II-B Remote Physiological Measurement
Recently, the combination of deep learning algorithms and rPPG technology has effectively improved accuracy in complex environments [50, 51]. Many studies have been conducted in the direction of multi-task monitoring of physiological data. For example, Liu et al. [52] presented a multi-task network to enable respiration and heart rate measurements on a mobile platform. Narayanswamy [53] et al. proposed a model based on facial movement, heart rate and respiration measurement. Wang et al. [27] designed a PhysMLE model to simultaneously calculate HR, heart rate variability, respiration rate, and blood oxygen saturation.
@@ II-C Non-contact Driver State Monitoring
Over the past few decades, with the development of deep learning technologies and increased demands for driving safety, non-contact driver state monitoring has become mainstream. Non-contact vision-based detection is categorized into conventional criterion-based methods and direct deep-learning detection methods. Specifically, using criteria combined with traditional methods such as eyes opening degree, mouth opening degree, or head posture, researchers were able to successfully detect driver states [54]. However, these criterion-based approaches are difficult to apply in diverse complex driving situations. Hence, direct fatigue detection algorithms based on deep learning algorithms have become a more feasible solution. For example, Ma et al. [55] proposed an image-based yawn detection algorithm that uses the driver’s image as an input for fatigue detection. However, this method can only detect fatigue by a single static image while losing temporal information in continuous video streams. Thus, video-based deep-learning models have also been proposed [56, 57, 58]. Although these methods are effective in obtaining time information, there are additional training and deployment costs. Recently, Mou et al. [19] developed a self-supervised learning algorithm using eye, nose, and head optical flow for fatigue detection. In addition, some studies [54, 59] used facial landmarks to detect driver drowsiness. For example, Yang et al. [17] used eye, mouth, and facial feature points as inputs to determine the level of driver fatigue.
However, to the best of our knowledge, no non-contact algorithms based on video images have been proposed for cognitive load assessment. Consequently, we introduce a novel non-contact driver monitoring system. Given that drowsiness and high cognitive load can both negatively affect driving performance in the context of driving automation, our method aims to estimate the driver’s drowsiness and cognitive load concurrently, utilizing video information. Additionally, it incorporates the remote Photoplethysmography (rPPG) technique for measuring physiological data, thereby enhancing the correlation between the various outputs of the model.
@@ III-A Problem Formulation
Inspired by this, we introduced a multi-task learning architecture, a video-based multi-task DMS mixture of experts model (VDMoE). As shown in Figure 1, suppose we have a batch of $B$ raw videos. Being different from previous methods taking 3D videos as input [56, 19], we extracted heterogeneous facial features from videos as input $X$. Firstly, following [17], we extracted facial landmarks, and subregion facial videos (i.e., eyes and mouse) from the video. Then, as only movement features from landmarks and above subregions might be insufficient for physiological measurement, we combined the knowledge from rPPG [60, 61] to supplement skin light changes into the input $X$. The specific preprocessing steps are in the following section. Finally, given $X$, we tried to obtain both physiological signals $\{h_{i},r_{i}\}_{i=1}^{N}$ and state estimation $\{d_{i},c_{i}\}_{i=1}^{N}$ by training a multi-task model $f(X;\theta)$, where $h_{i}$ is the heart rate, $r_{i}$ is the respiratory rate, $d_{i}$ is drowsiness levels and $c_{i}$ is the cognitive load levels. The trainable parameter $\theta$ is our optimization objective. Finally, the multi-task estimation target is $Y=\{h,r,d,c\}$
Figure 1: This figure presents the proposed multi-task driver state monitoring method VDMoE. We first transform videos into STMaps, facial landmarks matrix, and facial regions. Then inputs them into the model to output physiological signals, fatigue, and cognitive load.
@@ III-B Preprocess
At the beginning of preprocessing, we first applied the facial detection methods to each frame of the input video to generate 2D coordinates of $N_{f}$ facial landmarks with $N_{f}$ key points. Since the length of each video is $T$ frames and each facial landmark is represented by horizontal and vertical coordinates, denoted as $(x_{i},y_{i})$, each video can be represented as a facial landmarks matrix of shape $T\times N_{f}\times 2$. Then, by using the outer boundaries corresponding to the facial landmarks as bounding boxes, the regions of the left eye, right eye, and mouth $I_{l},I_{r},I_{m}$ were obtained from each original facial video frame. To standardize the input, the subregion video of the left eye and right eye $I_{l},I_{r}$ are resized to $T\times L_{e}\times W_{e}\times C$, where $L_{e},W_{e}$ is the length, width of subregion videos corresponding to eyes. Similarly, the subregion video of mouse $I_{m}$ is in the size of $T\times L_{m}\times W_{m}\times C$, where $L_{m},W_{m}$ is the length, and width of mouse video. Subsequently, the entire facial areas were cropped into $N_{s}$ regions of interest (ROIs) [62]. These ROIs were converted from RGB to YUV. Then, since we know that the skin light changes owing to the volume and oxygen saturation of the blood corresponds to the low amplitude of low-frequency component [63], we used the first-order difference to remove the large amplitude caused by head movements and change in light. After a Butterworth filter with parameter [0.4,10] Hz to retain the frequency band of interest [62], we processed the ROIs into a spatial-temporal map (STMap) $I_{s}\in\mathbb{R}^{T\times N_{s}\times C}$ to measure physiological signals. Among them, the channel number $C=3$ is constant. In all, the model’s input $X$ can be described as $X=\{I_{l},I_{r},I_{m},I_{f},I_{s}\}$.

### v2

@@ Abstract
Road safety remains a critical challenge worldwide, with approximately 1.35 million fatalities annually attributed to traffic accidents, often due to human error. As we advance towards higher levels of vehicle automation, challenges still exist, as driving with automation can cognitively over-demand drivers if engaging in non-driving-related tasks (NDRTs), or lead to drowsiness if driving is the sole task. This calls for the urgent need for an effective Driver Monitoring System (DMS) that can evaluate cognitive load and drowsiness in SAE Level-2/3 autonomous driving contexts. In this study, we propose a novel multi-task DMS, termed VDMoE, which leverages RGB video input to monitor driver states non-invasively. By utilizing key facial features to minimize computational load and integrating remote Photoplethysmography (rPPG) for physiological insights, our approach enhances detection accuracy while maintaining efficiency. Additionally, we optimize the Mixture-of-Experts (MoE) framework to accommodate multi-modal inputs and improve performance across different tasks. A novel prior-inclusive regularization method is introduced to align model outputs with statistical priors, thus accelerating convergence and mitigating overfitting risks. We validated our method by creating a new dataset (MCDD), which comprises RGB video and physiological indicators from 42 participants, as well as two public datasets. Our findings demonstrate the effectiveness of VDMoE in monitoring driver states, contributing to safer autonomous driving systems. The code and data will be released.
@@ I Introduction
Road safety is still an extreme challenge for societies [1, 2]. The World Health Organization (WHO) reported that approximately 1.35 million people die each year as a result of road traffic accidents, with human error being a significant contributing factor. While autonomous technologies promise to reduce human error—the leading cause of traffic accidents—there remains a critical period where human drivers must share control with automated systems. Particularly, at the Society of Automotive Engineers (SAE) Level-2 (L2) automation [3], the vehicle can control both steering and acceleration/deceleration but still requires the driver to remain actively engaged and ready to take over at any moment; while with SAE Level-3 (L3) automation, the vehicle can handle most driving tasks but still requires human drivers to step in when prompted by the take-over requests (TORs). Both have shifted the role of the driver from an operator to a supervisor. As a result, drivers are more likely to experience low cognitive load if driving is the sole task, or more likely to be cognitively over-demanding, if drivers are engaged in non-driving-related tasks (NDRTs), both can be detrimental to driving performance [4, 5]. For example, the cognitive overload induced by multi-dimensional cognitive resource consumption can lead to a limited field of vision [6], and a reduced ability to foresee potential hazards [7]. Further, the effect of long-time driving drowsiness and the high cognitive load as a result of NDRTs may co-exist, leading to impaired driving performance [8, 9]. Thus, the Driver Monitoring System (DMS) that is capable of estimating drivers’ cognitive load and drowsiness is critical to enhancing the safety and reliability of L2/L3 vehicles.
@@
However, most of the previous methods were based on single-frame detection, which loses temporal variation in both surface facial expression and internal physiological reflection. Compared to image-based methods, video-based methods [21, 18, 22] can make effective use of continuous temporal features and have the potential to enhance detection accuracy. However, existing methods still have the following shortcomings: (1) some video-based drowsiness detection methods [23] directly utilize continuous frames as the input, leading to significant computational resources and interference from redundant information; (2) some methods [18, 24] tried to replace video by key optimized facial features (e.g., landmarks, eye and mouth areas), while they cannot assess physiological features without supplementing signals from physiological sensors; (3) current video-based DMS focused mostly on single-task drowsiness [21, 18, 22] or distraction [25] estimation. However, driver states do not appear independently; instead, the interaction effects among multiple states make driver state estimation complicated. For example, in the context of conditional autonomous driving (i.e., driving with SAE L2/L3 vehicles), the cognitive load can moderate both drowsiness [26] and physiological reflection [27]. However, few non-contact solutions were proposed for cognitive load estimation, drowsiness detection and physiological reflections. Given that training separate models for each task incurs high deployment costs and reduces iterative efficiency [28], it is urgent to develop a unified model and make full use of the association across drivers’ multiple states.
@@
To address the aforementioned issues, we propose an RGB video-based multi-task DMS (VDMoE). Specifically, based on previous studies [18, 35], we first utilized the key facial features to reduce redundant information and computational costs. At the same time, considering cardiac and respiration activities are crucial predictors of driver states [27, 26], and are also important driver health indicators, we integrated the remote Photoplethysmography (rPPG) technology [36] into our model. Since the facial landmark and local eye and mouth region process contain limited temporal color change of the facial skin pixel points, which can be converted to a change of the blood volume under the skin [28], we first selected key facial regions of interest based on facial landmarks. Then, by transforming the color space (from RGB to YUV) as well as by band-pass denoising, we obtained the alternative multi-modal information input Spatial-temporal Map (STMap). Next, to fully leverage the dependencies between different tasks, the Mixture-of-Experts (MoE) [37] structure is introduced. Being different from the classic MoE structure, the heterogeneous gating mechanism and spatio-temporal expert separation are designed for the multi-modal alternative inputs to improve the multi-task performance. Besides, recent video-based DMS used various types of neural networks (e.g., CNN [21], Transformer [18]) that are suitable for processing different types of data or extracting different levels of feature information [38]. However, given the demand for real-time assessment of DMS, networks with excessive parameters and computational complexity will pose a challenge to deployment hardware. Therefore, we instantiated the nonlinear feature learning component of the basic block of VDMoE as a simple two-layer multi-layered perception (MLP) network to replace the previous modules represented by residual convolution [39] or multi-head attention [40]. Lastly, for the optimization goal, to explicitly capture the dependency between different states, we strengthen the learning capacity of VDMoE through one regularization based on prior knowledge from human factors fields [26]. Given there is no proper dataset for multi-task video-based DMS development that considered both the cognitive load and drowsiness in the driving automation context (seeing Table I), a driving simulator experiment with 42 participants was conducted. Extensive experiments based on our dataset illustrate the effectiveness of our method. In all, the contributions of our work are summarized as follows:
@@
  • A multi-task RGB video-based driver state monitoring method (VDMoE) is proposed in this work. This system concentrates on the drowsiness and physiological reflection moderated by the multi-dimensional cognitive load in the context of driving automation. As far as we know, this is the first multi-task model that can simultaneously estimate driver drowsiness, cognitive load, and cardiac and respiratory activities.
@@
  • To achieve the balance between estimation accuracy and efficiency, we replaced the full video with key facial features (landmarks, eye, and mouth area) as input, and introduced the STMap representing periodical color changes in facial regions to supplement alternative physiological features.
@@
  • Inspired by the superior performance of MoE in multi-task learning, we further optimized the MoE structure to accommodate the multi-modal and spatio-temporal input through a heterogeneous gating mechanism and spatio-temporal expert separation. At the same time, we replaced more complex network structures with MLPs for lightweight and future deployment. Comparison of computational cost with existing baselines proves our proposals’ effectiveness.
@@
  • A novel prior-inclusive regularization, inspired by human factor engineering studies, was proposed to enforce the multi-task model’s outputted probability distributions for each state to be aligned with the statistical prior distributions so that we can speed up convergence and reduce the risk of overfitting the model to particular individuals. Results from the ablation test validate its effectiveness.
@@
  • To validate the effectiveness of our method, the first multi-modal cognitive load and drowsiness driving dataset (MCDD) in the L3 context was created. In total, 42 subjects were involved in this dataset, which included about 105,840 seconds RGB video and cardiac and respiration-related physiological indicators. The cleaned processed dataset will be released.
@@ II-A Physiological Signal-Based Driver State Monitoring
Driver drowsiness [41] can negatively affect drivers’ performance when driving with automation, especially in takeover events. Thus, driver state monitoring is necessary to ensure safe driving in the context of driving automation [12]. According to [42, 43], in conditional autonomous vehicles, certain physiological metrics can be strongly correlated with driver state. Therefore, researchers proposed various physiological signal-based driver state monitoring methods. Commonly used psychophysiological measures include electrophysiological signals and human physical movements [44]. For example, Cui et al. [45] designed an interpretable convolutional neural network for drowsiness detection using EEG signals. In addition, due to the difficulty of obtaining EGG signals, methods based on Electrocardiography(ECG) signals [46] and Electrodermal Activity(EDA) signals [47] have also been widely proposed. Yang et al. [35] developed an attention-enabled recognition network with a decision-level fusion architecture to estimate cognitive load based on physiological data. Another study used physiological signals and facial images to detect driver drowsiness [21]. However, the above-mentioned studies used physiological data collected through intrusive sensors, i.e., the sensors had to contact with the driver’s body, which makes them infeasible to be used in the actual driving environment. Therefore, the VDMoE proposed in this paper uses a non-contact RGB camera as the perception method and simultaneously presents multiple driver states to replace the traditional driver monitoring based on contact sensors.
@@ II-B Remote Physiological Measurement
Recently, the combination of deep learning algorithms and rPPG technology has effectively improved accuracy in complex environments [51, 52]. Many studies have been conducted in the direction of multi-task monitoring of physiological data. For example, Liu et al. [53] presented a multi-task network to enable respiration and heart rate measurements on a mobile platform. Narayanswamy [54] et al. proposed a model based on facial movement, heart rate and respiration measurement. Wang et al. [28] designed a PhysMLE model to simultaneously calculate HR, heart rate variability, respiration rate, and blood oxygen saturation. Nevertheless, existing multi-task rPPG approaches typically focus on physiological signal estimation (e.g., HR and RR) or single psychological indicators, rather than jointly inferring both physiological and cognitive/drowsiness states. Consequently, they lack the robustness and comprehensive capability required for real-time driver monitoring under Level-2/3 automation, which our VDMoE model explicitly addresses.
@@ II-C Non-contact Driver State Monitoring
Over the past few decades, with the development of deep learning technologies and increased demands for driving safety, non-contact driver state monitoring has become mainstream. Non-contact vision-based detection is categorized into conventional criterion-based methods and direct deep-learning detection methods. Specifically, using criteria combined with traditional methods such as eyes opening degree, mouth opening degree, or head posture, researchers were able to successfully detect driver states [55]. However, these criterion-based approaches are difficult to apply in diverse complex driving situations. Hence, direct drowsiness detection algorithms based on deep learning algorithms have become a more feasible solution. For example, Ma et al. [56] proposed an image-based yawn detection algorithm that uses the driver’s image as an input for drowsiness detection. However, this method can only detect drowsiness by a single static image while losing temporal information in continuous video streams. Thus, video-based deep-learning models have also been proposed [57, 58, 59]. Although these methods are effective in obtaining time information, there are additional training and deployment costs. Recently, Mou et al. [20] developed a self-supervised learning algorithm using eye, nose, and head optical flow for drowsiness detection. In addition, some studies [55, 60] used facial landmarks to detect driver drowsiness. For example, Yang et al. [18] used eye, mouth, and facial feature points as inputs to determine the level of driver drowsiness.
While recent video-based DMS and rPPG research have begun exploring multi-task architectures, particularly those based on deep CNNs or Transformer backbones, these approaches still face substantial parameter overhead as tasks are added. Existing multi-task rPPG methods often suffer from model bloat due to task-specific decoupling layers and conflicting optimization objectives, leading to a ‘seesaw effect’ between tasks [28, 54]. Moreover, to our knowledge, no non-contact video-based algorithm has concurrently addressed cognitive load assessment, drowsiness detection, and physiological monitoring in Level-2/3 driving. In contrast, VDMoE leverages a MoE-like design that is inherently well-suited for multi-task learning: a small set of shared expert MLPs is dynamically routed by an element-wise gating network, and spatio-temporal feature extraction is separated across specialized experts. This architecture allows new tasks to be accommodated without a proportional increase in overall parameters. By adopting lightweight two-layer MLP experts instead of heavy CNN or Transformer modules [52, 53], VDMoE achieves notably smaller model size and lower inference latency while also integrating rPPG to capture physiological information. Consequently, VDMoE not only outperforms existing Transformer- or CNN-based multi-task models in efficiency and accuracy, but also uniquely provides a fully non-contact solution for simultaneous estimation of driver drowsiness, cognitive load, and physiological states in Level-2/3 automation contexts.
@@ III-A Problem Formulation
Inspired by this, we introduced a multi-task learning architecture, a video-based multi-task DMS mixture of experts model (VDMoE). As shown in Figure 3, suppose we have a batch of $B$ raw videos. Being different from previous methods taking 3D videos as input [57, 20], we extracted heterogeneous facial features from videos as input $X$. Firstly, following [18], we extracted facial landmarks, and subregion facial videos (i.e., eyes and mouth) from the video. Then, as only movement features from landmarks and above subregions might be insufficient for physiological measurement, we combined the knowledge from rPPG [61, 62] to supplement skin light changes into the input $X$. The specific preprocessing steps are in the following section. Finally, given $X$, we tried to obtain both physiological signals $\{h_{i},r_{i}\}_{i=1}^{N}$ and state estimation $\{d_{i},c_{i}\}_{i=1}^{N}$ by training a multi-task model $f(X;\theta)$, where $h_{i}$ is the heart rate, $r_{i}$ is the respiratory rate, $d_{i}$ is drowsiness levels and $c_{i}$ is the cognitive load levels. The trainable parameter $\theta$ is our optimization objective. Finally, the multi-task estimation target is $Y=\{h,r,d,c\}$
Figure 1: This figure presents the proposed multi-task driver state monitoring method VDMoE. We first transform videos into STMaps, facial landmarks matrix, and facial regions. Then inputs them into the model to output physiological signals, drowsiness, and cognitive load.
@@ III-B Preprocess
At the beginning of preprocessing, we first applied the facial detection methods to each frame of the input video to generate 2D coordinates of $N_{f}$ facial landmarks with $N_{f}$ key points. Since the length of each video is $T$ frames and each facial landmark is represented by horizontal and vertical coordinates, denoted as $(x_{i},y_{i})$, each video can be represented as a facial landmarks matrix of shape $T\times N_{f}\times 2$. Then, by using the outer boundaries corresponding to the facial landmarks as bounding boxes, the regions of the left eye, right eye, and mouth $I_{l},I_{r},I_{m}$ were obtained from each original facial video frame. To standardize the input, the subregion videos of the left eye and right eye $I_{l},I_{r}$ are resized to $T\times L_{e}\times W_{e}\times C$, where $L_{e},W_{e}$ are the length, width of the subregion videos corresponding to eyes. Similarly, the subregion video of mouth $I_{m}$ is the size of $T\times L_{m}\times W_{m}\times C$, where $L_{m},W_{m}$ are the length, and width of the mouth video. Subsequently, the entire facial areas were cropped into $N_{s}$ regions of interest (ROIs) [63]. These ROIs were converted from RGB to YUV. Then, since we know that the skin light changes owing to the volume and oxygen saturation of the blood correspond to the low amplitude of the low-frequency component [64], we used the first-order difference to remove the large amplitude caused by head movements and changes in light. After a Butterworth filter with parameter [0.4,10] Hz to retain the frequency band of interest [63], we processed the ROIs into a spatial-temporal map (STMap) $I_{s}\in\mathbb{R}^{T\times N_{s}\times C}$ to measure physiological signals. Among them, the channel number $C=3$ is constant. In all, the model’s input $X$ can be described as $X=\{I_{l},I_{r},I_{m},I_{f},I_{s}\}$.